OpenStudy (anonymous):

Dividing expressions. 8x^2/x+4 / x/2(x+5)

OpenStudy (anonymous):

\[\frac{\frac{8x^2}{x+4}}{\frac{x}{2(x+5)}}\]\[\frac{8x^2}{x+4} \div \frac{x}{2(x+5)}\]\[\frac{8x^2}{x+4} \times \frac{2(x+5)}{x}\]\[\frac{8x^2*2(x+5)}{(x+4)*x}\]\[\frac{16x^2(x+5)}{x(x+4)}\]\[\frac{16x(x+5)}{(x+4)}\]

OpenStudy (anonymous):

hey! thanks for answering :) can you explain the last two steps? please?

OpenStudy (anonymous):

since we have 16x^2/x, the x cancels out so we are left with 16x

OpenStudy (anonymous):

i'm guessing that's a rule?

OpenStudy (anonymous):

yes \[\frac{a^m}{a^n}=a^{m-n}\] :D

OpenStudy (anonymous):

ooooh okay :) thanks! :D